When Katy was 12, her father was three times her age. Now he is twice her age. How old is Katy now?
Answer by Edwin McCravy(20054)

When Katy was 12, her father was three times her age. Now he is twice her age. How old is Katy now?
Let Katy's age now = K
Let the Father's age now be F
Let x = the number of years ago it was when her father was 3 times her age.

When Katy was 12,
That was x years ago so
 
K - x = 12

her father was three times her age.
That was x years ago, so

F - x = 3(K - x)

Now he is twice her age
F = 2K

So we have this system of three equations:


K - x = 12
F - x = 3(K - x)
F = 2K

Simplifying them and putting the system in standard order

  K       -  x = 12
-3K +  F  + 2x =  0
-2K +  F       =  0

Solve that by elimination or substitution and get

K=24, F=48, x=12

Answer: Katy is 24, her father is 48, (48 is twice her age 24, that checks)
12 years ago Katy was 12 and her father was 36  (36 is 3 times 12,
that checks).

Edwin